OpenStudy (anonymous):

Find the circumference of the taxicab circle with a radius of 3. a. 24 c. 16 b. 12 d. 30 please help!

OpenStudy (anonymous):

im not really good at this but i would say ummm b 12

OpenStudy (sweetburger):

C=pi*2*r so just plug in the radius into this equation and you can find your circumference

OpenStudy (anonymous):

the answer is not shown because the answer is 18.84

OpenStudy (anonymous):

so i did 2*3=6 6*3.14=18.84

OpenStudy (anonymous):

no you have to do 2* 3.14* 3

OpenStudy (sweetburger):

well Im not sure where i went wrong tbh That is the circumference formula of a circle

OpenStudy (anonymous):

oh lol i said i wasnt good at these

OpenStudy (sweetburger):

@skylewis318 that is exactly what I said LOL

OpenStudy (anonymous):

i reallty good at theese

OpenStudy (anonymous):

does it want me to find the closest number?

OpenStudy (sweetburger):

@skylewis318 you do realize that 6x3.14 is the same thing as 2 x 3.14 x 3

OpenStudy (anonymous):

i do realize that now

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

would it be c?

OpenStudy (anonymous):

is c.16

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i would say tried but i wouldn't

OpenStudy (anonymous):

hmmm im confused because its not really giving me the right answer here even though it should be 18.84

OpenStudy (anonymous):

are u sure those are the answer choices

OpenStudy (anonymous):

yes im sure lol

OpenStudy (anonymous):

i would try 16

OpenStudy (anonymous):

can i have a medal

OpenStudy (anonymous):

i finially got it instead of muliplying the 3.14 in there you just do 2*3*4=24

OpenStudy (anonymous):

oh wow

OpenStudy (anonymous):

well i will see you later

OpenStudy (anonymous):

ok lol